(1-Dimensional Collision) A 100g solid ball approaches a 1kg solid ball which is initially at rest. If the initial speed of the 100g ball is 5m/s, what are the final speeds of the balls?


1
Expert's answer
2021-04-19T17:12:50-0400

Let a 100g solid ball have "m_A" and velocity "v_B"

Let a 1000g solid ball have "m_B" and velocity "v_A"

"m_Au_A+m_Bu_B=m_Av_A+m_Bv_B"

"0.5+0=0.1v_A+v_B"

"0.5_Au_A+0=0.1v_A+v_B"

"5=v_A+10v_B..........(1)"

If the collision is elastic (e=1)

"v_B-v_A=v_A-v_B"

"v_B-v_A=u_A"

"v_B-v_A=-5.........(2)"

Compairing the 2 equations

"v_A=v_B-5"

"5=-5 +v_B+10v_B"

"10=11v_B"

"v_B=0.909 m\/s"

"v_A=0.909-5=-4.091 m\/s"